OpenStudy (anonymous):

Find the probability for one roll of a number cube. P(3 or 4) 5/6, 2/3, 1/2, 1/6, 1/3 Urgent!

OpenStudy (amistre64):

how many numbers are on the cube?

OpenStudy (amistre64):

and, how many of them are a 3 or a 4?

OpenStudy (anonymous):

Those fractions are the numbers.

OpenStudy (amistre64):

and the solution will result when we can determine the answers to the questions i asked.

OpenStudy (amistre64):

we are not here to be a free answering service, but rather to try to help you learn the material.

OpenStudy (anonymous):

I know(: I literally just got on this subject today, I really need help on. I do not understand it at all.

OpenStudy (amistre64):

the basic mathematical definition of a probability is defined as:\[\frac{\#desired}{\#total}\] if we know the answers to the questions i asked then we could easily calculate the probability

OpenStudy (amistre64):

how many of them are a 3 or a 4? ----- divided by ------------- how many numbers are on the cube?

OpenStudy (amistre64):

for example: spose we had a ten sided dice, and six sides had a 3 or a 4 on them then the probability of rolling a 3 or a 4 would be: 6/10 = 3/5

OpenStudy (anonymous):

So basically the question is asking how likely is the dice to roll a 3 or 4?